### v1

@@ 5 Discussion
Relations between the de-reddened colours and the extinction-free colours Q, Q,Q,Q,..Q were constructed using a set of O-rich Mira-like stars with known interstellar extinction (Messineo et al. 2005). Such equations allow us to determine the interstellar extinction for every Mira-like star that has accessible near- and mid-infrared colors.
@@
In order to verify the applicability and reliability of the new methodology in determining interstellar extinction, an analysis of existing catalogs of Miras with available and mid-infrared measurements is made.
The most obscured Miras often exhibit OH maser emission and are therefore called OH/IR stars.
@@
Typically, their flux distribution is modeled with radiative transfer codes, and their interstellar extinction is derived from infrared extinction maps. For example, Olofsson et al. (2022) analyzed a sample of 22 AGB stars with OH masers near the Galactic center by using CO line observations taken with The Atacama large (sub-)millimeter array (Alma) array. The map of extinction by Gonzalez et al. (2018) and the extinction law by Nishiyama et al. (2009) were used. Archival photometry (from the 2MASS, VVV, and UKIDSS surveys 444VVV stands for the VISTA Variables in the Via Lactea survey (Soto et al. 2013). The UKIDSS Galactic Plane Survey (GPS) is described in Lucas et al. (2008)., the catalogues of Nogueras-Lara et al. (2019) and Nogueras-Lara et al. (2021)) and mid-infrared photometry from GLIMPSE surveys (5.8 and 8.0 um) could be retrieved from the VIZIER database for 8 stars of the Olofsson’s sample. The equations of Table 2 were used to make interstellar extinction estimates (A(int) from 0.90 to 2.0 mag). The two sets of A(int) linearly correlate with a standard deviation =0.49 mag and a mean difference of 0.09 mag (or 0.24 mag when rescaling the extinction to a power law with index = 2.1). The non-coevality of the near- and mid-infrared photometry accounts for the large ; the OH/IR stars have typical K amplitudes ranging from 0.9 to 2.5 mag (Messineo et al. 2020).
@@
Using data from the ISOGAL survey (Omont et al. 2003), Ojha et al. (2003) determined interstellar extinction, mass-loss rate, and luminosity values for 321 late-type stars located in Bulge fields at (, ). The majority of the sampled stars are AGB stars because they are brighter than the stars at the tip of the red giant branch. The authors used DENIS and ISOGAL photometry at 7 and 15 m and the silicate models by Jeong et al. (2003). The new method is used to estimate A(int) for 228 stars in this sample based on GLIMPSE and 2MASS data. The Ojha’s A values are converted to A using a factor of 0.089 (Ojha et al. 2003), and, then, compared with the A obtained here. The mean difference is 0.09 mag and a mag.
Matsunaga et al. (2009) estimate the A(int) values of 52 Mira stars at the Galactic center using average near-infrared measurements and the extinction law by Nishiyama et al. (2006).
@@
The average measurements along with the GLIMPSE [5.8] and [8.0] magnitudes are used to compute the A(int) values for 30 stars using the new method. The A(HK) from Matsunaga were converted to a power law with index = (by multiplying for 0.91) and range from 1.81 to 2.59 mag. An average difference of 0.1 mag is obtained and mag.
@@
The established equations between intrinsic infrared colours and extinction-free colours (Q), which are listed in Table 2, are useful to automatically determine stellar obscuration. This set of equations can be used to select late-type stars in a given sky region, and to improve their estimates of luminosity.
@@
Assuming a reasonable estimate of the total, the envelope extinction can be calculated as the difference between the total and the interstellar extinction. Mathematically, it is a straightforward calculation. In reality, estimating the circumstellar extinction is actually difficult because accurate spectral types are lacking, naked-star colors are unknown and affected by strong variable molecular absorption.
@@
Future photometric multi-wavelengths and multi-epochs surveys and spectroscopic surveys, such as LSST, Gaia DR4, 4MOST, GALAH, will enable us to obtain spectral types and temperatures, and their variations. This will allow us to perform more accurate extinction calculations.
Appendix A BC and intrinsic colours versus Q.
@@
In Messineo (2004), BC values as a function of infrared colours were obtained for the MSX filters. M values and dereddened WISE and GLIMPSE magnitudes from the revised catalog of Messineo et al. (2018) are here used to extend the BC computation to the WISE and Glimpse filters. Polynomial fits are listed in Table 8.
@@
X-axis | Y-axis | Coef n1 | Coef n2 | Coef n3 | Y-axis Y-Fit | X-range | |
(mag) | (mag) | ||||||
(KsA)o | BC(Ks) | 2.540 0.028 | 0.632 0.021 | 0.156 0.003 | 0.000 | 0.158 | |
(KsA)o | BC(A) | 2.540 0.028 | 1.632 0.021 | 0.156 0.003 | 0.000 | 0.158 | |
(KsC)o | BC(Ks) | 1.869 0.051 | 0.857 0.027 | 0.140 0.003 | 0.000 | 0.169 | |
(KsC)o | BC(C) | 1.869 0.051 | 1.857 0.027 | 0.140 0.003 | 0.000 | 0.169 | |
(KsD)o | BC(Ks) | 1.834 0.055 | 0.846 0.028 | 0.132 0.003 | 0.000 | 0.176 | |
(KsD)o | BC(D) | 1.834 0.055 | 1.846 0.028 | 0.132 0.003 | 0.000 | 0.176 | |
(KsE)o | BC(Ks) | 1.605 0.104 | 0.850 0.045 | 0.114 0.005 | 0.000 | 0.183 | |
(KsE)o | BC(E) | 1.605 0.104 | 1.850 0.045 | 0.114 0.005 | 0.000 | 0.183 | |
(KsW1)o | BC(Ks) | 3.088 0.026 | 0.154 0.037 | 0.239 0.011 | 0.000 | 0.248 | |
(KsW1)o | BC(W1) | 3.088 0.026 | 1.154 0.037 | 0.239 0.011 | 0.000 | 0.248 | |
(KsW2)o | BC(Ks) | 3.074 0.028 | 0.165 0.024 | 0.126 0.005 | 0.000 | 0.229 | |
(KsW2)o | BC(W2) | 3.074 0.028 | 1.165 0.024 | 0.126 0.005 | 0.000 | 0.229 | |
(KsW3)o | BC(Ks) | 2.459 0.047 | 0.562 0.027 | 0.117 0.004 | 0.000 | 0.218 | |
(KsW3)o | BC(W3) | 2.459 0.047 | 1.562 0.027 | 0.117 0.004 | 0.000 | 0.218 | |
(KsW4)o | BC(Ks) | 1.789 0.074 | 0.751 0.032 | 0.103 0.003 | 0.000 | 0.238 | |
(KsW4)o | BC(W4) | 1.789 0.074 | 1.751 0.032 | 0.103 0.003 | 0.000 | 0.238 | |
(Ks)o | BC(Ks) | 3.270 0.019 | 0.0790.022 | 0.162 0.009 | 0.000 | 0.265 | |
(Ks)o | BC( | 3.270 0.019 | 0.921 0.022 | 0.162 0.009 | 0.000 | 0.265 | |
(Ks)o | BC(Ks) | 3.158 0.021 | 0.199 0.032 | 0.177 0.010 | 0.000 | 0.243 | |
(Ks)o | BC() | 3.158 0.021 | 1.199 0.032 | 0.177 0.010 | 0.000 | 0.243 | |
(Ks)o | BC(Ks) | 2.884 0.027 | 0.462 0.027 | 0.162 0.006 | 0.000 | 0.198 | |
(Ks)o | BC() | 2.884 0.027 | 1.462 0.027 | 0.162 0.006 | 0.000 | 0.198 | |
(Ks)o | BC(Ks) | 2.781 0.034 | 0.506 0.028 | 0.141 0.005 | 0.000 | 0.201 | |
@@
The BC values and the intrinsic colours are plotted versus the interstellar-extinction-free parameter Qin Figs. 11, 12, and 13 for the MSX, WISE, and GLIMPSE filters.
Appendix B and the jolly equation
This is the definition of (Ks-) colour excess:
By combining Eq. B2. and B.3, it is obtained
and, therefore,
where
The B.5 is the definition of the parameter.
The first member contains observed quantities,
while the second term only contains intrinsic colours.
This equivalence proves that does not depend
on interstellar extinction. It is an intrinsic colour of the star.
Analogously, a function can be defined with the colour.

### v5

@@ 5 Discussion
Relations between the de-reddened colours and the extinction-free colours Q, Q,Q,QC,..QW4 were constructed using a set of O-rich Mira-like stars with known interstellar extinction (Messineo et al. 2005). Such equations allow us to determine the interstellar extinction for every Mira-like star that has accessible near- and mid-infrared (NIR-MIR) colours.
@@
In order to verify the applicability and reliability of the new methodology in determining interstellar extinction, we carried out an analysis of existing catalogs of Miras with available and MIR measurements.
@@
The most obscured Miras often exhibit OH maser emission and are therefore called OH/IR stars. Typically, their flux distribution is modeled with radiative transfer codes, and their interstellar extinction is derived from infrared extinction maps. For example, Olofsson et al. (2022) analyzed a sample of 22 AGB stars with OH masers near the Galactic center by using CO line observations taken with The Atacama large (sub-)millimeter array (Alma) array. The map of extinction by Gonzalez et al. (2018) and the extinction law by Nishiyama et al. (2009) were used. Archival photometry (from the 2MASS, VVV, and UKIDSS surveys666VVV stands for the VISTA Variables in the Via Lactea survey (Soto et al. 2013). The UKIDSS Galactic Plane Survey (GPS) is described in Lucas et al. (2008)., the catalogues of Nogueras-Lara et al. (2019) and Nogueras-Lara et al. (2021)) and mid-infrared photometry from GLIMPSE surveys (5.8 and 8.0 um) could be retrieved from the VIZIER database for 8 stars of the Olofsson’s sample. The equations of Table 2 were used to make interstellar extinction estimates (A(int) from 0.90 to 2.0 mag). The two sets of A(int) linearly correlate with a standard deviation =0.49 mag and a mean difference (this work - Olofsson’s values) of 0.09 mag (or 0.24 mag when rescaling the extinction to a power law with index = 2.1). The non-coevality of the near- and mid-infrared photometry accounts for the large ; the OH/IR stars have typical Ks amplitudes ranging from 0.9 to 2.5 mag (Messineo et al. 2020).
@@
Using data from the ISOGAL survey (Omont et al. 2003), Ojha et al. (2003) determined interstellar extinction, mass loss rate, and luminosity values for 321 late-type stars located in Bulge fields at (, ). The majority of the sampled stars are AGB stars because they are brighter than the stars at the tip of the red giant branch. The authors used DENIS and ISOGAL photometry at 7 and 15 m and the silicate models by Jeong et al. (2003). The new method was used to estimate AK(int) for 228 stars in this sample based on GLIMPSE and 2MASS data. The Ojha’s AV values are converted to A using a factor of 0.089 (Ojha et al. 2003) and then compared with the A obtained here. The mean difference is 0.09 mag and a mag.
@@
Matsunaga et al. (2009) estimate the AK(int) values of 52 Mira stars at the Galactic center using average near-infrared measurements and the extinction law by Nishiyama et al. (2006). The average measurements along with the GLIMPSE [5.8] and [8.0] magnitudes are used to compute the A(int) values for 30 stars using the new method. The A(HK) from Matsunaga were converted to a power law with index = (by multiplying for 0.91) and range from 1.81 to 2.59 mag. An average difference between Matsunaga’s extinction values and the here estimated values of +0.1 mag is obtained and mag.
@@
The established equations between intrinsic infrared colours and extinction-free colours (Qλ), which are listed in Table 2, are useful to automatically determine stellar obscuration. This set of equations can be used to select late-type stars in a given sky region, and to improve their estimates of luminosity.
@@
Assuming a reasonable estimate of the total, the envelope extinction can be calculated as the difference between the total and the interstellar extinction. Mathematically, it is a straightforward calculation. In reality, estimating the circumstellar extinction is actually difficult because accurate spectral types are lacking, naked-star colours are unknown and affected by strong variable molecular absorption.
@@
Future photometric multi-wavelengths and multi-epochs surveys and spectroscopic surveys, such as LSST, Gaia DR4, 4MOST, and GALAH, will enable us to obtain spectral types and temperatures, and their variations. As a result, we will be able to obtain more accurate extinction calculations.
Acknowledgements.
This work has made use of data from the European Space Agency (ESA) mission Gaia (), processed by the Gaia Data Processing and Analysis Consortium (DPAC, ). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center / California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. This work is based on observations made with the Spitzer Space Telescope, which is operated by the Jet Propulsion Laboratory, California Institute of Technology under a contract with NASA. This research made use of data products from the Midcourse Space Experiment, the processing of which was funded by the Ballistic Missile Defense Organization with additional support from the NASA office of Space Science. This publication makes use of data products from WISE, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory / California Institute of Technology, funded by the National Aeronautics and Space Administration. This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France, and SIMBAD database. This research utilized the NASA’s Astrophysics Data System Bibliographic Services. This work is based on the PhD thesis by Messineo M. (2004) which was supported by the Netherlands Research School for Astronomy (NOVA) through a netwerk 2, Ph.D. stipend. MM thanks Dr. Harm Habing and Dr. Frank Bertoldi for insightful discussions on interstellar extinction during her PhD thesis, and the anonymous referee for his constructive inputs.References
- Allard (2014) Allard, F. 2014, in Exploring the Formation and Evolution of Planetary Systems, ed. M. Booth, B. C. Matthews, & J. R. Graham, Vol. 299, 271–272
- Allard et al. (2011) Allard, F., Homeier, D., & Freytag, B. 2011, in Astronomical Society of the Pacific Conference Series, Vol. 448, 16th Cambridge Workshop on Cool Stars, Stellar Systems, and the Sun, ed. C. Johns-Krull, M. K. Browning, & A. A. West, 91
- Bayo et al. (2008) Bayo, A., Rodrigo, C., Barrado Y Navascués, D., et al. 2008, A&A, 492, 277
- Bedijn (1987) Bedijn, P. J. 1987, A&A, 186, 136
- Blommaert et al. (2018) Blommaert, J. A. D. L., Groenewegen, M. A. T., Justtanont, K., & Decin, L. 2018, MNRAS, 479, 3545
- Cami (2002) Cami, J. 2002, PhD thesis, University of Amsterdam, Netherlands
- Churchwell et al. (2009) Churchwell, E., Babler, B. L., Meade, M. R., et al. 2009, PASP, 121, 213
- Cohen et al. (2003) Cohen, M., Wheaton, W. A., & Megeath, S. T. 2003, AJ, 126, 1090
- David & Papoular (1990) David, P. & Papoular, R. 1990, A&A, 237, 425
- Draine (1985) Draine, B. T. 1985, ApJS, 57, 587
- Egan et al. (2003) Egan, M. P., Price, S. D., & Kraemer, K. E. 2003, in Air Force Research Laboratory Technical Report N. AFRL-VS-TR-2003-1589, Vol. 1, 1–89
- Epchtein et al. (1994) Epchtein, N., de Batz, B., Copet, E., et al. 1994, Ap&SS, 217, 3
- Eyer et al. (2023) Eyer, L., Audard, M., Holl, B., et al. 2023, A&A, 674, A13
- Fazio et al. (2004) Fazio, G. G., Hora, J. L., Allen, L. E., et al. 2004, ApJS, 154, 10
- Fritz et al. (2011) Fritz, T. K., Gillessen, S., Dodds-Eden, K., et al. 2011, ApJ, 737, 73
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Glass & Schultheis (2002) Glass, I. S. & Schultheis, M. 2002, MNRAS, 337, 519
- Gonzalez et al. (2018) Gonzalez, O. A., Minniti, D., Valenti, E., et al. 2018, MNRAS, 481, L130
- Gordon et al. (2021) Gordon, K. D., Misselt, K. A., Bouwman, J., et al. 2021, ApJ, 916, 33
- Habing et al. (2006) Habing, H. J., Sevenster, M. N., Messineo, M., van de Ven, G., & Kuijken, K. 2006, A&A, 458, 151
- Hauschildt et al. (1999) Hauschildt, P. H., Allard, F., Ferguson, J., Baron, E., & Alexander, D. R. 1999, ApJ, 525, 871
- Indebetouw et al. (2005) Indebetouw, R., Mathis, J. S., Babler, B. L., et al. 2005, ApJ, 619, 931
- Ivezic et al. (1999) Ivezic, Z., Nenkova, M., & Elitzur, M. 1999, arXiv e-prints, astro
- Jarrett et al. (2011) Jarrett, T. H., Cohen, M., Masci, F., et al. 2011, ApJ, 735, 112
- Jeong et al. (2003) Jeong, K. S., Winters, J. M., Le Bertre, T., & Sedlmayr, E. 2003, in Astrophysics and Space Science Library, Vol. 283, Mass-Losing Pulsating Stars and their Circumstellar Matter, ed. Y. Nakada, M. Honma, & M. Seki, 139–141
- Koornneef (1983) Koornneef, J. 1983, A&A, 128, 84
- Lebzelter et al. (2023) Lebzelter, T., Mowlavi, N., Lecoeur-Taibi, I., et al. 2023, A&A, 674, A15
- Lee et al. (2011) Lee, Y. S., Beers, T. C., An, D., et al. 2011, ApJ, 738, 187
- Lucas et al. (2008) Lucas, P. W., Hoare, M. G., Longmore, A., et al. 2008, MNRAS, 391, 136
- Maercker et al. (2022) Maercker, M., Khouri, T., Mecina, M., & De Beck, E. 2022, A&A, 663, A64
- Matsunaga et al. (2009) Matsunaga, N., Kawadu, T., Nishiyama, S., et al. 2009, MNRAS, 399, 1709
- Matsuura et al. (2002) Matsuura, M., Yamamura, I., Cami, J., Onaka, T., & Murakami, H. 2002, A&A, 383, 972
- Messineo (2004) Messineo, M. 2004, PhD thesis, Leiden Observatory, Leiden University, P.O. Box 9513, 2300 RA Leiden, The Netherlands
- Messineo (2022) Messineo, M. 2022, PASJ, 74, 1049
- Messineo et al. (2004) Messineo, M., Habing, H. J., Menten, K. M., Omont, A., & Sjouwerman, L. O. 2004, A&A, 418, 103
- Messineo et al. (2005) Messineo, M., Habing, H. J., Menten, K. M., et al. 2005, A&A, 435, 575
- Messineo et al. (2002) Messineo, M., Habing, H. J., Sjouwerman, L. O., Omont, A., & Menten, K. M. 2002, A&A, 393, 115
- Messineo et al. (2018) Messineo, M., Habing, H. J., Sjouwerman, L. O., Omont, A., & Menten, K. M. 2018, A&A, 619, A35
- Messineo et al. (2012) Messineo, M., Menten, K. M., Churchwell, E., & Habing, H. 2012, A&A, 537, A10
- Messineo et al. (2020) Messineo, M., Sjouwerman, L. O., Habing, H. J., & Omont, A. 2020, PASJ, 72, 63
- Nishiyama et al. (2006) Nishiyama, S., Nagata, T., Kusakabe, N., et al. 2006, ApJ, 638, 839
- Nishiyama et al. (2009) Nishiyama, S., Tamura, M., Hatano, H., et al. 2009, ApJ, 696, 1407
- Nogueras-Lara et al. (2019) Nogueras-Lara, F., Schödel, R., Gallego-Calvente, A. T., et al. 2019, A&A, 631, A20
- Nogueras-Lara et al. (2021) Nogueras-Lara, F., Schödel, R., & Neumayer, N. 2021, A&A, 653, A133
- Nogueras-Lara et al. (2020) Nogueras-Lara, F., Schödel, R., Neumayer, N., et al. 2020, A&A, 641, A141
- Ohnaka et al. (2016) Ohnaka, K., Weigelt, G., & Hofmann, K. H. 2016, A&A, 589, A91
- Ojha et al. (2003) Ojha, D. K., Omont, A., Schuller, F., et al. 2003, A&A, 403, 141
- Olofsson et al. (2022) Olofsson, H., Khouri, T., Sargent, B. A., et al. 2022, A&A, 665, A82
- Omont et al. (2003) Omont, A., Gilmore, G. F., Alard, C., et al. 2003, A&A, 403, 975
- Ossenkopf et al. (1992) Ossenkopf, V., Henning, T., & Mathis, J. S. 1992, A&A, 261, 567
- Perrin et al. (1998) Perrin, G., Coudé du Foresto, V., Ridgway, S. T., et al. 1998, A&A, 331, 619
- Price et al. (2001) Price, S. D., Egan, M. P., Carey, S. J., Mizuno, D. R., & Kuchar, T. A. 2001, AJ, 121, 2819
- Rieke & Lebofsky (1985) Rieke, G. H. & Lebofsky, M. J. 1985, ApJ, 288, 618
- Rosenthal et al. (2000) Rosenthal, D., Bertoldi, F., & Drapatz, S. 2000, A&A, 356, 705
- Scicluna et al. (2015) Scicluna, P., Siebenmorgen, R., Wesson, R., et al. 2015, A&A, 584, L10
- Skrutskie et al. (2006) Skrutskie, M. F., Cutri, R. M., Stiening, R., et al. 2006, AJ, 131, 1163
- Soto et al. (2013) Soto, M., Barbá, R., Gunthardt, G., et al. 2013, A&A, 552, A101
- Stead & Hoare (2009) Stead, J. J. & Hoare, M. G. 2009, MNRAS, 400, 731
- Suh (1999) Suh, K.-W. 1999, MNRAS, 304, 389
- Suh (2021) Suh, K.-W. 2021, ApJS, 256, 43
- van Belle et al. (2021) van Belle, G. T., von Braun, K., Ciardi, D. R., et al. 2021, ApJ, 922, 163
- van Loon et al. (2005) van Loon, J. T., Marshall, J. R., & Zijlstra, A. A. 2005, A&A, 442, 597
- Volk & Kwok (1988) Volk, K. & Kwok, S. 1988, ApJ, 331, 435
- Wang & Chen (2019) Wang, S. & Chen, X. 2019, ApJ, 877, 116
- Wang et al. (2021) Wang, T., Jiang, B., Ren, Y., Yang, M., & Li, J. 2021, ApJ, 912, 112
- Wright et al. (2010) Wright, E. L., Eisenhardt, P. R. M., Mainzer, A. K., et al. 2010, AJ, 140, 1868
Appendix A and the jolly equation
This is the definition of (Ks-) colour excess:
By combining Eq. A.2 and A.3, it is obtained
therefore,
where
The A.5 is the definition of the parameter. The first member contains observed quantities, while the second term only contains intrinsic colours. This equivalence proves that does not depend on interstellar extinction. It is an intrinsic colour of the star. Analogously, a function can be defined with the colour.
Appendix B BC and intrinsic colours versus Qλ.
@@ Acknowledgements.
In Messineo (2004), BC values as a function of infrared colours were obtained for the MSX filters. Mbol values and dereddened WISE and GLIMPSE magnitudes from the revised catalog of Messineo et al. (2018) are here used to extend the BC computation to the WISE and Glimpse filters. Polynomial fits are listed in Table 8.
@@
X-axis | Y-axis | Coef n1 | Coef n2 | Coef n3 | Y-axis Y-Fit | |
(mag) | (mag) | |||||
(KsA)o | BC(Ks) | 2.540 0.028 | 0.632 0.021 | 0.156 0.003 | 0.000 | 0.158 |
(KsA)o | BC(A) | 2.540 0.028 | 1.632 0.021 | 0.156 0.003 | 0.000 | 0.158 |
(KsC)o | BC(Ks) | 1.869 0.051 | 0.857 0.027 | 0.140 0.003 | 0.000 | 0.169 |
(KsC)o | BC(C) | 1.869 0.051 | 1.857 0.027 | 0.140 0.003 | 0.000 | 0.169 |
(KsD)o | BC(Ks) | 1.834 0.055 | 0.846 0.028 | 0.132 0.003 | 0.000 | 0.176 |
(KsD)o | BC(D) | 1.834 0.055 | 1.846 0.028 | 0.132 0.003 | 0.000 | 0.176 |
(KsE)o | BC(Ks) | 1.605 0.104 | 0.850 0.045 | 0.114 0.005 | 0.000 | 0.183 |
(KsE)o | BC(E) | 1.605 0.104 | 1.850 0.045 | 0.114 0.005 | 0.000 | 0.183 |
(KsW1)o | BC(Ks) | 3.088 0.026 | 0.154 0.037 | 0.239 0.011 | 0.000 | 0.248 |
(KsW1)o | BC(W1) | 3.088 0.026 | 1.154 0.037 | 0.239 0.011 | 0.000 | 0.248 |
(KsW2)o | BC(Ks) | 3.074 0.028 | 0.165 0.024 | 0.126 0.005 | 0.000 | 0.229 |
(KsW2)o | BC(W2) | 3.074 0.028 | 1.165 0.024 | 0.126 0.005 | 0.000 | 0.229 |
(KsW3)o | BC(Ks) | 2.459 0.047 | 0.562 0.027 | 0.117 0.004 | 0.000 | 0.218 |
(KsW3)o | BC(W3) | 2.459 0.047 | 1.562 0.027 | 0.117 0.004 | 0.000 | 0.218 |
(KsW4)o | BC(Ks) | 1.789 0.074 | 0.751 0.032 | 0.103 0.003 | 0.000 | 0.238 |
(KsW4)o | BC(W4) | 1.789 0.074 | 1.751 0.032 | 0.103 0.003 | 0.000 | 0.238 |
(Ks)o | BC(Ks) | 3.270 0.019 | 0.0790.022 | 0.162 0.009 | 0.000 | 0.265 |
(Ks)o | BC( | 3.270 0.019 | 0.921 0.022 | 0.162 0.009 | 0.000 | 0.265 |
(Ks)o | BC(Ks) | 3.158 0.021 | 0.199 0.032 | 0.177 0.010 | 0.000 | 0.243 |
(Ks)o | BC() | 3.158 0.021 | 1.199 0.032 | 0.177 0.010 | 0.000 | 0.243 |
(Ks)o | BC(Ks) | 2.884 0.027 | 0.462 0.027 | 0.162 0.006 | 0.000 | 0.198 |
(Ks)o | BC() | 2.884 0.027 | 1.462 0.027 | 0.162 0.006 | 0.000 | 0.198 |
(Ks)o | BC(Ks) | 2.781 0.034 | 0.506 0.028 | 0.141 0.005 | 0.000 | 0.201 |
@@
The BC values and the intrinsic colours are plotted versus the interstellar-extinction-free parameter Qλ in Figs. 11, 12, and 13 for the MSX, WISE, and GLIMPSE filters.